OpenStudy (sparklestaraa):

The graph below shows three functions:

OpenStudy (sparklestaraa):

Which function has all real numbers as its domain? Only f(x) Both g(x) and p(x) Both f(x) and g(x) Only p(x)

OpenStudy (sparklestaraa):

@bibby

OpenStudy (alexandervonhumboldt2):

fx only

OpenStudy (danjs):

The blue and the red functions have limited intervals, they are line segments

OpenStudy (danjs):

p(x), looks like the function y=e^x, and has a domain of all real numbers, even though it looks like it only goes to around x = 2, if you use HUGE y values, x does have all possible values, all real numbers

OpenStudy (danjs):

For that second one, first let x =0, and a number to the zero power is 1

OpenStudy (sparklestaraa):

ok :D

OpenStudy (danjs):

Then use f(1) and see what that point is, or use the fact that it DECAYS over time, so it is the one of the two left that is decreasing

OpenStudy (danjs):

F(x) is not the right answer for the first one, as alex said, dont know if you caught that

OpenStudy (sparklestaraa):

oh so 1st is p(x) ?

OpenStudy (sparklestaraa):

c?

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (sparklestaraa):

@AlexandervonHumboldt2 is it line D for the second and p(x) for the first?

OpenStudy (alexandervonhumboldt2):

for secnod D

OpenStudy (alexandervonhumboldt2):

first px

OpenStudy (danjs):

the second is not D, wrong again mr alex

OpenStudy (danjs):

f(0) = 4*3^0 = 4*1 = 4

OpenStudy (danjs):

and it is decreasing

OpenStudy (sparklestaraa):

B then?